OpenStudy (dragon123mama):

what is the expression in factored form? 2x^2 + 16x +24

OpenStudy (mariamikayla):

2x^2+16x+24= 2(x^2+8x+12)= =2(x^2+6x+2x+12)= =2( x(x+6)+2(x+6))= =2(x+6)(x+2)

OpenStudy (dragon123mama):

ohh i see. i wasn't sure how to solve this. thank you.